A system recieves 100 calories and the work done by the system is 418 joules.what is the change in its internal energy?
1
Expert's answer
2017-05-09T10:06:09-0400
100 calories = 418.4 joules
The internal energy is calculated as
U=Q+W, where Q is the hear received and W is the work done over the gas.
In our problem
U=Q-W= 418.4-418=0.4 J